OpenStudy (anonymous):

find the next two terms in the sequence 1,4,9

OpenStudy (anonymous):

16,25

OpenStudy (anonymous):

16,25 increasing with the next even number

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 1^2=1}\) \(\normalsize\color{blue}{ 2^2=4}\) \(\normalsize\color{blue}{ 3^2=9}\) \(\normalsize\color{blue}{ 4^2=16}\) \(\normalsize\color{blue}{ 5^2=25}\) \(\normalsize\color{blue}{ 6^2=36}\) and so on ....

OpenStudy (solomonzelman):

All perfect squares in a row.